# ClaimDB: A Fact Verification Benchmark over Large Structured Data

### v1

@@ Abstract
Despite substantial progress in fact-verification benchmarks, claims grounded in large-scale structured data remain underexplored. In this work, we introduce ClaimDB, the first fact-verification benchmark where the evidence for claims is derived from compositions of millions of records and multiple tables. ClaimDB consists of 80 unique real-life databases covering a wide range of domains, from governance and healthcare to media, education and the natural sciences. At this scale, verification approaches that rely on “reading” the evidence break down, forcing a timely shift toward reasoning in executable programs. We conduct extensive experiments with 30 state-of-the-art proprietary and open-source (below 70B) LLMs and find that none exceed 83% accuracy, with more than half below 55%. Our analysis also reveals that both closed- and open-source models struggle with abstention—the ability to admit that there is no evidence to decide—raising doubts about their reliability in high-stakes data analysis. We release the benchmark, code, and the LLM leaderboard at https://claimdb.github.io.
@@ 1 Introduction
In this work, we propose ClaimDB, a fact-verification benchmark in which claims are grounded in evidence deliberately composed from multiple tables and millions of rows. Successful verification in ClaimDB implicitly requires some form of executable reasoning  Chen et al. (2023) as the evidence is too large to naively “read”, consume and combine. Each claim is paired with a unique database (out of 80) that contains, on average, 11 tables and 4.5M records, mirroring the scale and complexity of real-world fact-checking.
@@ 2 Related Work
Fact verification started with free-form text. The seminal work of Thorne et al. (2018) on FEVER pairs claims with Wikipedia sentences to determine whether they are supported or refuted. LIAR Wang (2017) is a complementary benchmark of political claims with fine-grained veracity labels. Since then, fact-verification has evolved beyond text to include structured and semi-structured evidence. FEVEROUS Aly et al. (2021) extends FEVER with tables, while other works explore verification over time-series data Strong and Vlachos (2025), Wikipedia info-boxes Gupta et al. (2020), knowledge graphs Dammu et al. (2024), financial reports Zhao et al. (2024), and temporal data Barik et al. (2024, 2025).
@@
However, existing fact-verification benchmarks rely on tables that are small enough to fit within an LLM’s context window. Evidence comes almost exclusively from Wikipedia or scientific articles, where tables are inherently small. In ClaimDB we break this pattern by designing a benchmark where claims are grounded in databases with millions of records. At this scale, verification can no longer rely on naively “reading” and then reasoning over in-context evidence.
@@
Property Value
Total Claims 53,368
Label distribution
# Entailed (E) 12,855
# Contradicted (C) 16,529
# Not Enough Info (NEI) 23,984
   NEI subcategories
   # Out-of-Schema 12,644
   # Counterfactual 5,786
   # Subjective 5,554
Claim Context
Databases per Claim 1
Avg. Tables per Database 11.6
Avg. Records per Database 4.6M
Total Databases (DBs) 80
Table 1: Summary statistics of ClaimDB.
3 ClaimDB Benchmark
In ClaimDB, the task is to determine whether a given claim is 1 entailed, 2 contradicted, or 3 has not-enough-info (NEI) with respect to the evidence in a database. Each claim is paired with exactly one large, multi-table database, which serves as the sole source of evidence for verification. The workflow for creating the ClaimDB benchmark is shown in Figure 1. In this section we describe the first half of the workflow—the creation of all claims (Section 4 covers the quality-control half).
@@ 3.1 Overview
What makes ClaimDB challenging is the scale of the underlying data. Each database contains, on average, 11.6 tables and 4.6M records. At this scale, common approaches that rely on feeding tables directly into an LLM (e.g., Wang et al., 2024c) are no longer viable. Instead, successful systems must rely on executable programs to carry out the heavy compositional reasoning required by our benchmark—for example, averaging or sorting over millions of records (Section 3.3). Promising solutions include but are not limited to PoT Chen et al. (2023) and tool-calling agents Theologitis and Suciu (2025).
@@ 3.2 BIRD Benchmark
BIRD provides us with 11k NL-to-SQL examples (the combined public train and test splits). Each example is paired with the underlying database and accompanied by external information specific to each example. For example, in the toxicology database, a label marked with “++” indicates that a molecule is carcinogenic, while “-” indicates the opposite. Throughout all subsequent transformations, we preserve any external info intact. Importantly for us, the NL-to-SQL pairs have been carefully curated and vetted by trained, independent annotators, and the SQL queries are produced via a double-blind annotation process with experts in the loop.
@@ 3.3 Pre-Filtering
The common theme is that answering the question (or, later, verifying the claim) requires combining information from large portions of the database, which will far exceed the context size of any LLM.
Finally, we discard queries whose answers return more than ten records. Answers are fed to an LLM during claim generation (Section 3.5) so this prevents overwhelming the model with bookkeeping.
@@ 3.5 Claim Generation
Our final setup uses a single prompt for each label (3 in total). All final prompts are available in Figures 12, 13 and 14 of the Appendix.
For both entailed and contradicted claims, we found that 1-shot prompting with medium reasoning works best. Given a single Q/A pair, the model generates between one and three claims.
@@
First, we generate claims that fall outside the concepts of the database. To do this, we inspect the database schema metadata offline—including table and column names, and relationships—and provide the full schema information to gpt-5. The model has no trouble using this effectively: for example, in Figure 3, the generated out-of-schema claim refers to homeschooling, which is not represented anywhere in the California Schools database.
@@
For NEI claim generation, we found that zero-shot prompting with reasoning works best. Any form of in-context learning consistently hurt generation quality: gpt-5 is already capable of producing these claims reliably, and few-shot prompting mainly ends up constraining its imagination.
Finally, we end up with roughly 14k, 19k, and 28k entailed, contradicted, and NEI claims, respectively (Figure 1). The discrepancy in counts arises because contradicted claims are not constrained by the answer itself (they can oppose it arbitrarily), while NEI claims are not constrained by either the answer or even the concepts present in the Q/A pair or the database (we address this in Section 4.2).
@@ 4 Quality Control
Automatic claim generation is never perfect. Issues can still slip-in such as claims that reference prior context—which is opaque to a verifier—or leak helpful information. While outright label errors are rare (the claim generation task is trivial), we take conservative steps to filter problematic claims. This is the second half of the workflow illustrated in Figure 1.
@@ 4.1 Panel of LLM Judges
Prompt Engineering. After calibration, we end up with the prompts shown in Figures 15 and 16 of the Appendix. The former is used for entailed and contradicted claims, while the latter for NEI claims and includes the two additional evaluation rubrics.
Figure 4: Performance of the judge panel on the human-annotated test set (75 NEI and 75 C or E). The panel achieves $100\%$ recall on the examples that the human annotators found problematic and behaves conservatively.
Prompt Results. The judge panel achieves $100\%$ recall on the claims flagged by human annotators; they are all eliminated by the panel as well. We also did not observe any mislabeled claims. This aligns with our experience during the construction of the benchmark—generating claims from correct Q/A pairs is an easy task, and gpt-5 is more than capable of doing so reliably. Figure 4 shows the behavior of the panel on the annotated test set.
@@ 4.2 Grounding NEI Claims
The main reason we have a large number of NEI claims is that they are inherently unconstrained by the database concepts. They can be about almost anything, especially in the case of out-of-schema claims. However, this freedom often makes them very obvious to classify. The same issue arises with counterfactual claims—given a bit of imagination (gpt-5 does not lack this quality), it is easy to go overboard. These overboard claims are not bad per se; we simply want a way to avoid sampling them when we construct our main benchmarking test set.
@@
Figure 5: Distribution of smilarity scores between generated claims and their gold context for out-of-schema claims. Higher scores indicate that claims stay “closer” to the underlying data concepts. The two highlighted claim examples ($\clubsuit$, $\bigstar$) are discussed in Section 4.2.
@@ 5 Benchmark Splits
We divide ClaimDB into three disjoint splits: 1 the training set, 2 a public test set, and 3 a private test set (the labels are hidden). Both test sets contain approximately 1,000 examples each. They are sampled [^2] The test sets are drawn from claims coming from the dev split of BIRD as it has received substantial cleanup effort in prior work (e.g., Liu et al., 2025; Wretblad et al., 2024). uniformly at random per label and fully balanced, with roughly 333 examples per label. For NEI claims, the three subcategories are also balanced with roughly 111 examples each; for both out-of-schema and counterfactual, sampling is further restricted to the top quartile of the similarity score distribution (Figures 5 and 9).
@@ 6.1 Setup
Claims in ClaimDB are intentionally derived from insights that combine information across millions of records through operations such as averaging, grouping and sorting (Section 3.3). At this scale, verification cannot be done by “reading” tables and reasoning in-context: the databases contain 4.6M records on average—which is also far beyond any LLM’s context window. As a result, verifiers must reason in executable programs that handle the bulk of the compositional reasoning that is required.
Our evaluation primarily focuses on tool-calling agents Wang et al. (2024a); Mialon et al. (2024), which we view as a very strong and practical solution for this setting. We use Google’s MCP toolbox for databases Buvaraghan and Egan (2025) and give each agent a single tool: the ability to execute arbitrary SQL queries against the database associated with the claim at hand. The only constraint we impose is a generous limit of 20 tool calls; exceeding this counts as a failed run.
@@ 6.3 Discussion
We analyze results on the public test set. Results on the private test set are reported in Appendix A.2.
LLM Backbone Precision Recall F1 Macro-F1 Acc. Total Tokens
@@
gpt-4o-mini (2024) 0.777 0.666 0.679 0.577 0.656 0.107 0.662 0.661 0.185 0.503 0.445 27M
gpt-4.1-nano (2025a) 0.647 0.641 0.376 0.165 0.124 0.952 0.263 0.208 0.539 0.337 0.416 13M
gpt-5-nano (2025b) 0.815 0.710 0.874 0.742 0.900 0.720 0.777 0.794 0.790 0.787 0.787 24M
gpt-5-mini (2025b) 0.875 0.713 0.963 0.754 0.952 0.777 0.810 0.815 0.860 0.828 0.827 14M
gpt-oss:20b (2025) 0.851 0.735 0.676 0.669 0.688 0.862 0.749 0.710 0.758 0.739 0.740 14M
gemini-2.5-flash (2025) 0.841 0.734 0.830 0.685 0.825 0.869 0.755 0.777 0.849 0.793 0.793 6M
gemini-3-flash (2025) 0.754 0.750 0.962 0.799 0.934 0.673 0.776 0.832 0.792 0.800 0.801 10M
claude-3-haiku (2024) 0.572 0.556 0.433 0.261 0.435 0.756 0.359 0.488 0.550 0.466 0.485 9M
claude-3-5-haiku (2024) 0.775 0.626 0.661 0.580 0.743 0.702 0.663 0.680 0.681 0.675 0.675 11M
claude-haiku-4-5 (2025) 0.836 0.711 0.983 0.796 0.952 0.682 0.815 0.814 0.805 0.811 0.809 38M
qwen3:1.7b (2025) 0.645 0.533 0.351 0.060 0.048 0.982 0.110 0.089 0.518 0.239 0.366 2M
qwen3:4b (2025) 0.530 0.554 0.492 0.450 0.218 0.860 0.487 0.312 0.626 0.475 0.511 11M
qwen3:8b (2025) 0.785 0.617 0.450 0.306 0.311 0.940 0.441 0.414 0.608 0.488 0.521 10M
qwen3:14b (2025) 0.812 0.744 0.438 0.285 0.360 0.943 0.422 0.485 0.599 0.502 0.531 9M
qwen3:32b (2025) 0.732 0.667 0.491 0.393 0.459 0.866 0.512 0.544 0.626 0.561 0.574 10M
qwen3-coder:30b (2025) 0.744 0.625 0.660 0.646 0.659 0.711 0.691 0.641 0.685 0.672 0.672 22M
qwq:32b (2025) 0.500 0.558 0.472 0.336 0.508 0.667 0.402 0.532 0.552 0.495 0.504 9M
mistral-nemo:12b (2024) 0.380 0.446 0.381 0.381 0.211 0.577 0.381 0.287 0.459 0.376 0.391 8M
ministral-3:3b (2025b) 0.511 0.500 0.390 0.276 0.163 0.821 0.359 0.246 0.529 0.378 0.422 16M
ministral-3:8b (2025b) 0.633 0.557 0.484 0.408 0.396 0.792 0.496 0.463 0.600 0.520 0.533 17M
ministral-3:14b (2025b) 0.612 0.638 0.624 0.598 0.580 0.690 0.605 0.608 0.655 0.623 0.623 22M
mistral-small:22b (2025c) 0.423 0.526 0.375 0.222 0.060 0.878 0.291 0.108 0.525 0.308 0.389 16M
magistral:24b (2025a) 0.682 0.714 0.419 0.309 0.257 0.905 0.426 0.378 0.573 0.459 0.492 18M
devstral:24b (2025a) 0.415 0.624 0.399 0.291 0.221 0.771 0.342 0.326 0.526 0.398 0.429 19M
devstral-small-2 (2025b) 0.602 0.712 0.521 0.610 0.447 0.705 0.606 0.549 0.599 0.585 0.588 15M
nemotron-3-nano (2025) 0.714 0.726 0.612 0.652 0.601 0.747 0.681 0.658 0.673 0.671 0.667 29M
llama3.1:8b (2024) 0.341 0.433 0.337 0.222 0.079 0.726 0.269 0.133 0.461 0.288 0.344 7M
llama3.2:3b (2024) 0.428 0.412 0.366 0.372 0.100 0.685 0.398 0.161 0.477 0.345 0.387 8M
cogito:14b (2025) 0.758 0.701 0.384 0.141 0.184 0.973 0.238 0.292 0.551 0.360 0.435 8M
cogito:32b (2025) 0.797 0.564 0.506 0.354 0.556 0.792 0.491 0.560 0.617 0.556 0.568 11M
@@
Table 2: Results on the public test set. We report per-label precision, recall, and F1, along with macro-F1 and accuracy. We also report the total number of tokens (input + output) across the 1,000 experiments for each model.
Figure 6: Final accuracy versus model size.
@@
Scaling Trends. Figure 6 plots accuracy against model size. We also see the familiar pattern of larger models performing better, but with improvements that are marginal and roughly log-linear.
@@
Long sessions hurt. Long LLM-database interactions have inherent fragility: as they grow longer, models increasingly lose focus due to the large amounts of information. In this setting, a single bad decision—a single careless query—can flood the model with hundreds of thousands of values.
Very Expensive. To meaningfully fact-check a claim, a verifier must truly understand the data at its disposal. This inevitably requires consuming—either intentionally or unintentionally—a lot of information. As shown in Table 2, models consume and produce on the order of tens of millions of tokens over the full test set. For example, running claude-haiku-4-5 (the cheapest model in the 4.5 series) cost us \$43 for the public test split alone. This highlights the need to balance performance with cost in real-life settings like ClaimDB.

### v2

@@ Abstract
Real-world fact-checking often involves verifying claims grounded in structured data at scale. Despite substantial progress in fact-verification benchmarks, this setting remains largely underexplored. In this work, we introduce ClaimDB, a fact-verification benchmark where the evidence for claims is derived from compositions of millions of records and multiple tables. ClaimDB consists of 80 unique real-life databases covering a wide range of domains, from governance and healthcare to media, education and the natural sciences. At this scale, verification approaches that rely on “reading” the evidence break down, forcing a timely shift toward reasoning in executable programs. We conduct extensive experiments with 30 state-of-the-art proprietary and open-source (below 70B) LLMs and find that more than half score below 55% accuracy. Our analysis also reveals that both closed- and open-source models struggle with abstention—the ability to admit that there is no evidence to decide—raising doubts about their reliability in high-stakes data analysis tasks. We release the benchmark, code, and the LLM leaderboard at https://claimdb.github.io.
@@ 1 Introduction
In this work, we propose ClaimDB, a fact-verification benchmark in which claims are grounded in evidence deliberately composed from multiple tables and millions of rows. Successful verification in ClaimDB implicitly requires some form of executable reasoning  Chen et al. (2023) as the evidence is too large to naively “read”, consume, and combine. Each claim is paired with a unique database—out of 80 in total—that corresponds to an evidence context of roughly 110M tokens. More specifically, each claim involves 11 tables and 4.5M records on average, reflecting the scale and complexity of real-world fact-checking.
@@
The remainder of the paper is organized as follows. Section 2 reviews existing fact-checking benchmarks. Sections 35 describe the full benchmark construction pipeline, including claim generation, quality control with LLM judges, and dataset splits. We then evaluate 30 LLMs under a specific neuro-symbolic architecture in Section 6, before concluding the paper in Section 7.
@@ 2 Related Work
Fact verification started with free-form text. The seminal work of Thorne et al. (2018) on FEVER pairs claims with Wikipedia sentences to determine whether they are supported or refuted. LIAR Wang (2017) is a complementary benchmark of political claims with fine-grained veracity labels. Since then, fact-verification has evolved beyond text to include structured and semi-structured evidence. FEVEROUS Aly et al. (2021) extends FEVER with tables, while Bussotti et al. (2023, 2024) automate training-data generation for such settings. Other works explore verification over time-series data Strong and Vlachos (2025), Wikipedia info-boxes Gupta et al. (2020), knowledge graphs Dammu et al. (2024), financial reports Zhao et al. (2024), temporal data Barik et al. (2024, 2025), scientific paper abstracts Wadden et al. (2020), U.S. court rulings Putta et al. (2026), and combined text and visual representations of tabular data (e.g., charts) Zhou et al. (2025).
@@
However, existing fact-verification benchmarks rely on tables that are small enough to fit within an LLM’s context window. Evidence is drawn almost exclusively from Wikipedia or scientific articles, where tables are inherently limited in size. Recent work by Devasier et al. (2025, 2026) begins to explore larger structured data, but these approaches remain limited—either to small-scale pilot claim sets or to single-table evidence. In ClaimDB, we go further by designing a benchmark where claims are grounded in databases with millions of records, requiring multi-hop and compositional reasoning across tables. At this scale and complexity, verification can no longer rely on naively “reading” and then reasoning over in-context evidence.
3 ClaimDB Benchmark
In ClaimDB, the task is to determine whether a given claim is 1 entailed, 2 contradicted, or 3 has not-enough-info (NEI) with respect to the evidence in a database. Each claim is paired with exactly one large, multi-table database, which serves as the sole source of evidence for verification.
The workflow for creating the ClaimDB benchmark is shown in Figure 1. In this section, we focus on the first stage of this process—the creation of claims—followed by quality-control in Section 4.
Property Value
Total Claims 53,368
Label distribution
# Entailed (E) 12,855
# Contradicted (C) 16,529
# Not Enough Info (NEI) 23,984
   NEI subcategories
   # Out-of-Schema 12,644
   # Counterfactual 5,786
   # Subjective 5,554
Claim Context (Evidence)
Databases per Claim 1
Avg. Tables per Claim 11.3
Avg. Records per Claim 4.6M
Avg. Tokens per Claim 110M
Total Databases (DBs) 80
Table 1: Summary statistics of ClaimDB. Token counts per claim are computed using the tokenizer of the gpt-5 series as reference (see Appendix A for details).
@@ 3.1 Overview
What makes ClaimDB challenging is the scale of the underlying data. Each claim is paired with a database containing, on average, 11.3 tables and 4.6M records. When converted to text, this amounts to roughly 110M tokens of evidence per claim (Appendix A). This exceeds the context window of modern LLMs by two to three orders of magnitude—for reference, Google’s flagship long-context models support up to 1M–2M tokens Doshi (2025). Thus, common approaches that rely on feeding tables directly into an LLM (e.g., Wang et al., 2024c) are no longer viable.
Instead, successful systems must rely on executable programs to carry out the heavy compositional reasoning required by ClaimDB over large volumes of data—for example, averaging or sorting millions of records (Section 3.3). Promising directions include Program of Thoughts (PoT) Chen et al. (2023), tool-calling agents Wang et al. (2024a); Theologitis and Suciu (2025), and Recursive Language Models Zhang et al. (2026).
@@ 3.2 BIRD Benchmark
BIRD provides us with 11k NL-to-SQL examples (the combined public train and test splits). Each example is paired with the underlying database and accompanied by external information specific to each example. For example, in the toxicology database, a label marked with “++” indicates that a molecule is carcinogenic, while “-” indicates the opposite. Throughout all subsequent transformations, we preserve any external info intact. Importantly for us, the NL-to-SQL pairs have been carefully curated and vetted by trained, independent annotators, and the SQL queries are produced via a double-blind annotation process with experts in the loop Li et al. (2023).
@@ 3.3 Pre-Filtering
The common theme is that answering the question (or, later, verifying the claim) requires combining information from large portions of the database, which will far exceed the context size of any LLM. Figures 1316 in Appendix D provide concrete examples of queries, their ASTs, and the corresponding filtering decisions following the rules above.
Finally, we discard queries whose answers return more than ten records. These answers are later fed to an LLM during claim generation (Section 3.5), and large result sets would make it harder for the model to track the structure (columns and values). This ten-record cutoff is conservative for modern models like gpt-5, but helps ensure that errors do not arise from misinterpreting the structured data.
@@ 3.5 Claim Generation
Our final setup uses a single prompt for each label (3 in total). All final prompts are available in Figures 17, 18 and 19 of the Appendix.
For both entailed and contradicted claims, we found that 1-shot prompting with medium reasoning works best. Given a Q/A pair, the model is instructed to generate between one and three claims.
@@
First, we generate claims that fall outside the concepts of the database. To do this, we inspect the database schema metadata offline—including table and column names, and relationships—and provide the full schema information to gpt-5. We then explicitly instruct the model (Figure 19 in the Appendix) to produce claims that are based on concepts not represented anywhere in the schema. The model has no trouble doing this effectively: for example, in Figure 3, the generated out-of-schema claim refers to homeschooling—a concept that is not present in the California Schools database.
@@
For NEI claim generation, we found that zero-shot prompting with medium reasoning works best. There is a large spectrum of possible and appropriate NEI claims, as they are largely unconstrained by the concepts of the data. As a result, any form of in-context learning biases the model toward specific patterns, limiting generation diversity overall. In contrast, gpt-5 naturally generates a diverse set of NEI claims in the zero-shot setting.
All in all, we end up with roughly 14k, 19k, and 28k entailed, contradicted, and NEI claims, respectively (Table 1). The discrepancy in counts arises because contradicted claims are not constrained by the answer itself (they can oppose it arbitrarily), while NEI claims are not constrained by either the answer or even the concepts present in the Q/A pair or the database (we address this in Section 4.2).
@@ 4 Quality Control
Automatic claim generation is never perfect. Issues can still slip-in such as claims that reference prior context—which is opaque to a verifier—or leak helpful information. While outright label errors are rare (the claim generation task is trivial), we take conservative steps to filter problematic claims. This is the second half of the workflow in Figure 1.
@@ 4.1 Panel of LLM Judges
Prompt Engineering. After calibration, we end up with the prompts shown in Figures 20 and 21 of the Appendix. The former is used for entailed and contradicted claims, while the latter for NEI claims and includes the two additional evaluation rubrics.
Prompt Results. We evaluate the judge panel on the human-annotated test set (75 NEI and 75 contradicted or entailed), which contains 4 claims identified as problematic by PhD annotators. The panel successfully discards all four, achieving $100\%$ recall, and behaves conservatively. We also did not observe any mislabeled claims in terms of veracity in this set. This aligns with our experience during the construction of the benchmark—generating claims from correct Q/A pairs is an easy task, and gpt-5 is more than capable of doing so reliably.
@@ 4.2 Grounding NEI Claims
The main reason we have a large number of NEI claims is that they are inherently unconstrained by the database concepts. They can be about almost anything, especially in the case of out-of-schema claims. However, this freedom often makes them very obvious to classify. The same issue arises with counterfactual claims—given a bit of imagination (gpt-5 does not lack this quality), it is easy to go overboard. These overboard claims are not bad per se; we simply want a way to avoid sampling them when we construct our main benchmarking test set. In this section, we create a way to filter out claims based on how far they are from the underlying database concepts. We later use this signal to construct a more difficult test set (Section 5).
@@
Figure 4: Distribution of similarity scores between generated claims and their gold context for out-of-schema claims. Higher scores indicate that claims stay “closer” to the underlying data concepts. The two highlighted claim examples ($\clubsuit$, $\bigstar$) are discussed in Section 4.2.
@@ 5 Benchmark Splits
We divide ClaimDB into three disjoint splits: 1 the training set, 2 a public test set, and 3 a private test set (the labels are hidden). Both test sets contain approximately 1,000 examples each. They are sampled [^2] The test sets are drawn from claims coming from the dev split of BIRD as it has received substantial cleanup effort in prior work (e.g., Wretblad et al., 2024; Liu et al., 2025). uniformly at random per label and fully balanced, with roughly 333 examples per label. For NEI claims, the three subcategories are also balanced with roughly 111 examples each; for both out-of-schema and counterfactual, sampling is further restricted to the top quartile of the similarity score distribution (Figures 4 and 10). In contrast, the training split contains the full spectrum of NEI claims. We release similarity scores for all NEI claims, enabling the construction of any custom downstream training split—for example, one that mirrors the test set distribution.
@@ 6.1 Setup
Claims in ClaimDB are intentionally derived from insights that combine information across millions of records through operations such as averaging, grouping and sorting (Section 3.3). At this scale, verification cannot be done by “reading” the evidence in-context: the claim context contains, on average, 110M tokens of structured information—which is far beyond any LLM’s context window. As a result, verifiers must reason in executable programs that handle, symbolically, the bulk of the compositional reasoning that is required.
Our evaluation primarily focuses on tool-calling agents Mialon et al. (2024); Wang et al. (2024a), which we view as a very strong and practical solution for this setting. We use Google’s MCP toolbox for databases Buvaraghan and Egan (2025) and give each agent a single tool: the ability to execute arbitrary SQL queries against the database associated with the claim at hand. The only constraint we impose is a generous limit of 20 tool calls; exceeding this counts as a failed run.
@@ 6.3 Discussion
LLM Backbone Precision Recall F1 Macro-F1 Acc.
@@
gpt-4o-mini (2024) 0.777 0.666 0.679 0.577 0.656 0.107 0.662 0.661 0.185 0.503 0.445
gpt-4.1-nano (2025a) 0.647 0.641 0.376 0.165 0.124 0.952 0.263 0.208 0.539 0.337 0.416
gpt-5-nano (2025b) 0.815 0.710 0.874 0.742 0.900 0.720 0.777 0.794 0.790 0.787 0.787
gpt-5-mini (2025b) 0.875 0.713 0.963 0.754 0.952 0.777 0.810 0.815 0.860 0.828 0.827
gpt-oss:20b (2025) 0.851 0.735 0.676 0.669 0.688 0.862 0.749 0.710 0.758 0.739 0.740
gemini-2.5-flash (2025) 0.841 0.734 0.830 0.685 0.825 0.869 0.755 0.777 0.849 0.793 0.793
gemini-3-flash (2025) 0.754 0.750 0.962 0.799 0.934 0.673 0.776 0.832 0.792 0.800 0.801
claude-3-haiku (2024) 0.572 0.556 0.433 0.261 0.435 0.756 0.359 0.488 0.550 0.466 0.485
claude-3-5-haiku (2024) 0.775 0.626 0.661 0.580 0.743 0.702 0.663 0.680 0.681 0.675 0.675
claude-haiku-4-5 (2025) 0.836 0.711 0.983 0.796 0.952 0.682 0.815 0.814 0.805 0.811 0.809
qwen3:1.7b (2025) 0.645 0.533 0.351 0.060 0.048 0.982 0.110 0.089 0.518 0.239 0.366
qwen3:4b (2025) 0.530 0.554 0.492 0.450 0.218 0.860 0.487 0.312 0.626 0.475 0.511
qwen3:8b (2025) 0.785 0.617 0.450 0.306 0.311 0.940 0.441 0.414 0.608 0.488 0.521
qwen3:14b (2025) 0.812 0.744 0.438 0.285 0.360 0.943 0.422 0.485 0.599 0.502 0.531
qwen3:32b (2025) 0.732 0.667 0.491 0.393 0.459 0.866 0.512 0.544 0.626 0.561 0.574
qwen3-coder:30b (2025) 0.744 0.625 0.660 0.646 0.659 0.711 0.691 0.641 0.685 0.672 0.672
qwq:32b (2025) 0.500 0.558 0.472 0.336 0.508 0.667 0.402 0.532 0.552 0.495 0.504
mistral-nemo:12b (2024) 0.380 0.446 0.381 0.381 0.211 0.577 0.381 0.287 0.459 0.376 0.391
ministral-3:3b (2025b) 0.511 0.500 0.390 0.276 0.163 0.821 0.359 0.246 0.529 0.378 0.422
ministral-3:8b (2025b) 0.633 0.557 0.484 0.408 0.396 0.792 0.496 0.463 0.600 0.520 0.533
ministral-3:14b (2025b) 0.612 0.638 0.624 0.598 0.580 0.690 0.605 0.608 0.655 0.623 0.623
mistral-small:22b (2025c) 0.423 0.526 0.375 0.222 0.060 0.878 0.291 0.108 0.525 0.308 0.389
magistral:24b (2025a) 0.682 0.714 0.419 0.309 0.257 0.905 0.426 0.378 0.573 0.459 0.492
devstral:24b (2025a) 0.415 0.624 0.399 0.291 0.221 0.771 0.342 0.326 0.526 0.398 0.429
devstral-small-2 (2025b) 0.602 0.712 0.521 0.610 0.447 0.705 0.606 0.549 0.599 0.585 0.588
nemotron-3-nano:30b (2025) 0.714 0.726 0.612 0.652 0.601 0.747 0.681 0.658 0.673 0.671 0.667
llama3.1:8b (2024) 0.341 0.433 0.337 0.222 0.079 0.726 0.269 0.133 0.461 0.288 0.344
llama3.2:3b (2024) 0.428 0.412 0.366 0.372 0.100 0.685 0.398 0.161 0.477 0.345 0.387
cogito:14b (2025) 0.758 0.701 0.384 0.141 0.184 0.973 0.238 0.292 0.551 0.360 0.435
cogito:32b (2025) 0.797 0.564 0.506 0.354 0.556 0.792 0.491 0.560 0.617 0.556 0.568
@@
Table 2: We report per-label precision, recall, and F1, along with macro-F1 and accuracy.
Figure 5: Final accuracy vs. model size (open-source).
@@
Scaling Trends. Figure 5 plots accuracy against model size. We notice the familiar pattern of larger models performing better, but with improvements that are marginal and roughly log-linear.
Long Sessions Hurt. Long LLM-database interactions have inherent fragility: as they grow longer, models increasingly lose focus due to the large amounts of information. In this setting, a single bad decision—a single careless query—can flood the model with hundreds of thousands of tokens.
@@
Figure 7: Our agents are equipped with a single tool: the ability to execute SQL queries on the associated database. We report the success rates of the SQL queries per model. More specifically, “Success” denotes queries that execute and return results; “SQL Error” denotes queries with syntax or logical issues; and “Other Error” includes all remaining failures. Notably, top-performing models achieve high query success rates, suggesting that failures are not superficial (e.g., syntax) but instead reflect deeper breakdowns in reasoning over large structured data.
No Evidence of Data Contamination. A common concern with modern benchmarks is the risk of data contamination, where models may have seen parts of the data during training, making the benchmark obsolete Samuel et al. (2025). To investigate this, we evaluate the best-performing model, gpt-5-mini, without access to external tools, where it must label each claim using only its parametric knowledge. We find that performance is near random, with a macro-F1 of 0.253 and accuracy of 0.367, suggesting that ClaimDB cannot be solved using parametric knowledge alone.
@@
Notably, the performance gap between closed- and open-source models is driven mostly by how NEI is handled. However, neither behavior is desirable: proprietary models effectively avoid abstention, while open-source models over-abstein.
Non-Superficial Failures. Our neuro-symbolic baseline operates as a multi-step process driven by SQL queries over the underlying database. Hence, inspecting these queries provides insight into where model failures occur. In Figure 7, we plot the distribution of whether executed SQL queries succeed or fail. The two best-performing models, gpt-5-mini and claude-haiku-4.5, achieve SQL query success rates of 93% and 99%, respectively. For both models, the remaining failed queries are evenly distributed across correct and incorrect predictions. This suggests that, for stronger models, failures are not superficial (e.g., syntax), but instead reflect breakdowns in reasoning—specifically, in how the model composes queries to navigate the database.